OpenStudy (anonymous):

Solve by factoring. 3x2−30x=−75

OpenStudy (anonymous):

please help someone

OpenStudy (raden):

first, move sides -75 to LHS, be 3x^2−30x+75=0

OpenStudy (anonymous):

shouldnt i divide by 3 first to make the fraction problemeasier to work with: x2-10x+25=0

OpenStudy (anonymous):

?

OpenStudy (campbell_st):

divide each term by 3 x^2 - 10x + 25 = 0 looks like a perfect square

OpenStudy (anonymous):

@campbell_st thats what i got but then i get stuck because its te;;ing me my answers are wrong

OpenStudy (anonymous):

ll*

OpenStudy (anonymous):

x=5 PROOF: 3x^2-30x=-75 transpose 75 to the other side of the equation makes the equation 3x^2-30x+75 divide the equation by 3 makes it x^2 - 30x+25=0 factor the equation (x-5)^2 therefore x=5 checking: 3(5^2)-30(5)=-75 3(25)-150=-75 75-150=-75 -75=-75 therefore, 5 satisfies the equation

OpenStudy (raden):

x^2 - 10x + 25 = 0 (x-5)(x-5) = 0

OpenStudy (anonymous):

ohhh ok i see:)